Some long-range navigation systems use hyperbolas to determine a ship’s position. Suppose the system
imposes coordinates so that the location of a ship is in the first quadrant. A ship is located at the intersection of the hyperbolas with equations 25x^2 – 4y^2 = 100 and  36y^2 – x^2= 64 . Find the coordinates of the ship to the nearest hundredth of a unit.
a. (1.90, 4.28)
b. (2.07, 1.38)
c. (1.38, 2.07)

Answer by Theo(13342)

the solution is selection b. (2.07,1.38)

you can solve it algebraically as follows:

start with:

25x^2 – 4y^2 = 100
36y^2 – x^2 = 64

re-arrange the terms so that they line up like for like.

25x^2 - 4y^2 = 100
-x^2 + 36y^2 = 64

multiply both sides of the second equation by 25 to get:

25x^2 - 4y^2 = 100
-25x^2 + 900y^2 = 1600

add the equations together to get:

896y^2 = 1700

solve for y^2 to get:

y^2 = 1.897321429

solve for y to get:

y = plus or minus 1.377432913

since the intersection point has to be in the first quadrant, then you get:

y = plus 1.377432913

replace y^2 with 1.897321429 in the first equation to get:

25x^2 - 4 * 1.897321429 = 100

simplify this to get:

25x^2 - 7.589285714 = 100

add 7.589285714 to both sides of the equation to get:

25x^2 = 107.5892857

solve for x^2 to get:

x^2 = 107.5892857/25 = 4.303571429

solve for x to get:

x = plus or minus 2.074505104.

since x has to be in the first quadrant, then you get:

x = plus 2.074505104.

your intersection point in the first quadrant is:

(x,y) = (2.074505104,1.377432913)

round this to 2 decimal digits and you get:

(x,y) = (2.07,1.38)

that's selection b.

graphically, this looks like this:

$$$